OpenStudy (anonymous):

Find the slope of the tangent line to the graph of the function at the given point. g(x): (2/3)x^2 - (x/6) at (-1, (5/6) )

OpenStudy (anonymous):

g'(x) = (-4/3)x + 5/6 = (-4/3)(-1) + 5/6 = 13/6 = 2.17

OpenStudy (anonymous):

\[g'(x) = (4/3)x - (1/6)\] \[g'(-1) = -(4/3) - (1/6)\] get your denominator common and you'll find it to be -(3/2) as your final answer. Remember, derivatives ARE slopes of tangent lines!